Consider the sample 14, 19, 18, 20, 16, 9, 10, 6, and 8. Find the mean.


1
Expert's answer
2022-06-02T12:38:59-0400

We have sample values 14, 19, 18, 20, 16, 9, 10, 6,8, sample size N=9.



"\bar{x}=\dfrac{1}{9}(14+19+18+20+16+9""+10+6+8)=\dfrac{40}{3}"